if the height and area of a triangle increase by 20% and 8%respectively , find the % change in the base of the triangle?
Answer by Alan3354(31507)

b & h = base and height before the increase
B & H = after
-----------
area = bh/2
1.08*bh/2 = BH/2 = B*h*1.2/2
0.54bh = 0.6*h*B
54b = 60B
B = b*(54/60)
B = 0.9b
Change = 10% decrease